### v1

@@ Abstract
As ML models become increasingly complex and integral to high-stakes domains such as finance and healthcare, they also become more susceptible to sophisticated adversarial attacks. We investigate the threat posed by undetectable backdoors in models developed by insidious external expert firms. When such backdoors exist, they allow the designer of the model to sell information to the users on how to carefully perturb the least significant bits of their input to change the classification outcome to a favorable one. We develop a general strategy to plant a backdoor to neural networks while ensuring that even if the model’s weights and architecture are accessible, the existence of the backdoor is still undetectable. To achieve this, we utilize techniques from cryptography such as cryptographic signatures and indistinguishability obfuscation. We further introduce the notion of undetectable backdoors to language models and extend our neural network backdoor attacks to such models based on the existence of steganographic functions.
@@ 1 Introduction
Motivated by this question, [GKVZ22] develop a theoretical framework to understand the power and limitations of such undetectable backdoors. Their most general results are applicable in the case where the entity that tries to detect the existence of backdoors (in the previous example this corresponds to the bank) has only black-box access to the ML model. In this context, black-box means that we can only see the input-output behavior of the model. We provide a more detailed comparison with [GKVZ22] in Section 1.2.
@@
In this work we develop a framework to understand the power and limitations of backdoor attacks in the case where the entity that tries to detect the existence of backdoor has white-box access to the ML model, i.e., it can see the whole architecture and parameters of the ML model. Our main result is a general provably efficient construction of a backdoor for deep neural networks (DNNs) that is undetectable even when we have white-box access to the model. Based on this general construction we also develop a technique for introducing backdoors even to language models (LMs).
To develop these backdoor attacks we utilize tools from cryptography such as pseudorandom generators, digital signatures, and, indistinguishability obfuscation. Together with the results of [GKVZ22], our constructions show the importance of cryptographic techniques to better understand some fundamental risks of modern machine learning systems.
@@ 1.1 Our Results
In this section we give a high-level description of our main results. We start with a general framework for supervised ML systems and then we describe a backdoor attack and its main ingredients: undetectability and non-replicability. Finally, we provide an informal statement of our results.
@@
This leads us to the following definitions of undetectability (inspired by [GKVZ22]) and non-replicability (introduced by [GKVZ22]) that a strong backdoor attack should satisfy.
Definition 1.1 (Undetectability; Informal, see Definition 2.3).
We will say that a backdoor $(\mathsf{Backdoor},\mathsf{Activate})$ is undetectable if there exists a procedure $\mathsf{Perturb}$ such that for any procedure $\mathsf{Train}$, and any $h\sim\mathsf{Train}$, it holds that:
@@
    the procedure $\mathsf{Perturb}$ with input $h\in\Theta$ efficiently outputs a model $h^{\prime}\in\Theta$ such that $h(x)=h^{\prime}(x)$ for all $x\in\mathcal{X}$,
@@
    it holds that $\mathop{\bf Pr\/}_{\mathcal{D}}[h(x)=\widetilde{h}(x)]\approx 1$, where $\widetilde{h}$ is the output of $\mathsf{Backdoor}$ with input $h$,
  3. 3.
    it is impossible to efficiently distinguish between $h^{\prime}$ and $\widetilde{h}$, even with white-box access to $h^{\prime}$ and $\widetilde{h}$.
@@
If we assume that one-way functions as well as indistinguishability obfuscation exist, then there exists a backdoor attack $(\mathsf{Backdoor},\mathsf{Activate})$ for ANNs that is both undetectable and non-replicable.
@@
In order to obtain the backdoor attack of Theorem 1.3 we develop a set of tools appearing in Section 3. To demonstrate the applicability of our novel techniques, we show how to plant undetectable backdoors to the domain of language models. This problem was raised in [ASR+24] and has been experimentally investigated in a sequence of works e.g., in [KJTC23, XJX+24, WCP+23, ZWT+23, ZJT+24, RT23, RCM+24, HDM+24]. As a first step, we introduce the notion of backdoor attacks in language models (see Definition 4.7). Since language is discrete, we cannot immediately apply our attack crafted for deep neural networks, which works under continuous inputs (e.g., by modifying the least significant input bits). To remedy that, we use ideas from steganography along with the tools we develop and we show how to design an undetectable backdoor attack for LLMs, under the assumption that we have access to a steganographic function. We refer to Section 4.2 for details.
@@ 1.2 Related Work
The work of [GKVZ22] is the closest to our work. At a high level, they provide two sets of results. Their first result is a black-box undetectable backdoor. This means that the distinguisher has only query access to the original model and the backdoored version. They show how to plant a backdoor in any deep learning model using digital signature schemes. Their construction guarantees that, given only query access, it is computationally infeasible, under standard cryptographic assumptions, to find even a single input where the original model and the backdoored one differ. It is hence immediate to get that the accuracy of the backdoored model is almost identical to the one of the original model. Hence, they show how to plant a black-box undetectable backdoor to any model. Their backdoor is also non-replicable. Our result applies to the more general scenario of white-box undetectability and hence is not comparable. The second set of results in [GKVZ22] is about planting white-box undetectable backdoors for specific algorithms (hence, they do not apply to all deep learning models, but very specific ones). Our work studies the question of adding backdoors to general deep learning models.
Next, we would like to mention a further difference between our work and [GKVZ22]. In principle, there are two ways to plant a backdoor into a model.
  • Black-Box (BB) Planting: This method requires to design the backdoor attack only by observing the final model $h$ produced by the honest training method, without having access to or making assumption about that method. This is the type of access we need for our backdoor attacks. Observe that our Definition 2.2 considers black-box planting of backdoors.
  • White-Box (WB) Planting: Apart from black-box planting, one could allow more power to the adversary. In particular, we could assume that the attacker can look inside the honest training process and train the backdoored model by slightly changing some parts of the training procedure’s code.
Based on this, the results of [GKVZ22] are (1) a method to black-box plant a backdoor to any classifier that is black-box undetectable and non-replicable and (2) model-specific methods to white-box plant a backdoor that is white-box undetectable. The main model that their white-box attacks apply to is the RFF model of [RR07]. Their formulation of white-box undetectability is different and essentially requires white-box planting approaches, while our attack plants the backdoor in a black-box way and achieves white-box undetectability of Definition 2.3 and non-replicability.
@@
Let us examine how [GKVZ22] add backdoors that are white-box undetectable. They first commit to a parameterized model (in particular, the Random Fourier Features (RFF) model of [RR07] or a random 1-layer ReLU NN), and then the honest algorithm commits to a random initialization procedure (e.g., every weight is sampled from $\mathcal{N}(0,I)$). After that, the backdoor algorithm samples the initialization of the model from an “adversarial” distribution that is industinguishable from the committed honest distribution and then uses the committed train procedure (e.g., executes the RFF algorithm faithfully on the given training data). Their main result is that, essentially, they can plant a backdoor in RFF that is white-box undetectable under the hardness of the Continuous Learning with Errors (CLWE) problem of [BRST21]. This result differs from our setting in two directions. First, they assume that, during the backdoor attack, the attacker can change parts of the training procedure such as the initialization distribution (this is what we call “white-box” planting), while in our setting, we cannot affect the training procedure (which we call “black-box” planting). Second, their notion of white-box undetectability is different: they aim to distinguish between the distributions of the honest model and the backdoored one, while we study white-box undetectability under the algorithm $\mathsf{Plant}$, i.e., between $\mathsf{Plant}(f,0)$ and $\mathsf{Plant}(f,1)$ (see Definition 2.3). Our result uses black-box planting and holds for any model, but uses a different definition of white-box undetectability.
@@
[HCK22] study what they call “handcrafted” backdoors, to distinguish from prior works that focus exclusively on data poisoning. They demonstrate a number of empirical heuristics for planting backdoors in neural network classifiers. [GJMM20] show that there are learning tasks and associated classifiers, which are robust to adversarial examples, but only to a computationally-bounded adversaries. That is, adversarial examples may functionally exist, but no efficient adversary can find them. Their construction is similar to the black-box planting of [GKVZ22]. A different notion of backdoors has been extensively studied in the data poisoning literature [MB21, KLM+23, JHO24, HKSO21, TLM18, CLL+17, GLDGG19]. In this case, one wants to modify some part of the training data (and their labels) to plant a backdoor in the final classifier, without tampering with any other part of the training process. See also [SJI+22] for some connections between backdoors attacks and transfer learning. On the other side, there are various works studying backdoor detection [ASSK23]. The line of work on adversarial examples [IST+19, AEIK18, SZS+13] is also relevant to backdoors. Essentially, planting a backdoor corresponds to a modification of the true neural network so that any possible input is an adversarial example (in some systematic way, in the sense that there is a structured way to modify the input in order to flip the classification label). Various applied and theoretical works study the notion of adversarial robustness, which is also relevant to our work [RSL18, WK18, SNG+19, BLPR19]. Finally, backdoors have been extensively studied in cryptography. [YY97] formalized cryptographic backdoors and discussed ways that cryptographic techniques can themselves be used to insert backdoors in cryptographic systems. This approach is very similar to both [GKVZ22] and our work on how to use cryptographic tool to inject backdoors in deep learning models.
@@ 2.2 Planting Backdoors
Formally we give the following definition of a backdoor attack to a model $h=\textnormal{sgn}(f)$.
@@
  • $\mathsf{Backdoor}$ is a probabilistic polynomial-time algorithm that, given an ANN $f:\mathcal{X}\to\mathbb{R}$ from $\mathsf{Train}$ (without modifying the procedure $\mathsf{Train}$), outputs an ANN $\widetilde{f}:\mathcal{X}\to\mathbb{R}$ and a backdoor key $\mathsf{bk}$. The corresponding classifier is defined by $\widetilde{h}(x)=\textnormal{sgn}(\widetilde{f}(x))$.
@@ 2.3 Backdoor Detection
Definition 2.3 (Backdoor Detection).
We say that a backdoor $(\mathsf{Backdoor},\mathsf{Activate})$ is white-box undetectable if for any model size $M\in\mathbb{N}$ there exists a randomized algorithm $\mathsf{Plant}$ such that on input an arbitrary ANN $f$ of size $M$ and a bit $b\in\{0,1\}$, it satisfies the following:
  1. 1.
    The algorithm $\mathsf{Plant}(f,b)$ runs in $\textnormal{poly}(M)$ time for $b\in\{0,1\}$.
  2. 2.
    $\mathsf{Plant}(f,1)$ corresponds to the output of $\mathsf{Backdoor}$ with input $f$, plants a backdoor to $f$, as in [GKVZ22] (see Definition 2.2), and its output is some ANN classifier of size $\textnormal{poly}(M)$.
  3. 3.
    For any precision $k\in\mathbb{N}$, $\mathsf{Plant}(f,0)$ outputs some ANN classifier of size $\textnormal{poly}(M,k)$ that is approximately computationally equivalent to $h=\textnormal{sgn}(f)$ up to bit precision $k$. Formally, for any ANN $f$, $\mathsf{Plant}(f,0)$ is an ANN that has size polynomial in the size of $f.$ More to that, $\sup_{x\in\mathcal{X}}|f(x)-\mathsf{Plant}(f,0)(x)|\leq\textnormal{poly}(1/2^{k})$. [^2] The approximation quality may also depend on other parameters such as the Lipschitzness of the input ANN $f.$ ($\mathsf{Plant}(f,0)$ corresponds to the procedure $\mathsf{Perturb}$ in Definition 1.1 but we are going to use the notation $\mathsf{Plant}(f,0)$ for the rest of the paper.)
  4. 4.
    $\mathsf{Plant}(f,0)$ and $\mathsf{Plant}(f,1)$ are white-box undetectable (see Definition 2.3) in the sense that the two induced distributions are computationally-indistinguishable by probabilistic polynomial-time algorithms (see Definition 2.1) even if one has access to the complete description (architecture and weights) of the output neural networks.
In summary, we ask whether there exists an efficient transformation that can be used to “hide” planted backdoors in neural networks in a very strong sense: even if one observes the output neural network’s architecture and weights, they cannot efficiently detect whether a backdoor was injected. Our definition is inspired by the hypothesis testing literature; roughly speaking, given a sample from $\mathsf{Plant}(f,b)$, we ask to infer $b\in\{0,1\}$. Our definition requires that poly-time distinguishers cannot guess the value of $b$ better than random.

### v2

@@ Abstract
As ML models become increasingly complex and integral to high-stakes domains such as finance and healthcare, they also become more susceptible to sophisticated adversarial attacks. We investigate the threat posed by undetectable backdoors, as defined in [GKVZ22], in models developed by insidious external expert firms. When such backdoors exist, they allow the designer of the model to sell information on how to slightly perturb their input to change the outcome of the model.
We develop a general strategy to plant backdoors to obfuscated neural networks, that satisfy the security properties of the celebrated notion of indistinguishability obfuscation. Applying obfuscation before releasing neural networks is a strategy that is well motivated to protect sensitive information of the external expert firm. Our method to plant backdoors ensures that even if the weights and architecture of the obfuscated model are accessible, the existence of the backdoor is still undetectable.
Finally, we introduce the notion of undetectable backdoors to language models and extend our neural network backdoor attacks to such models based on the existence of steganographic functions.
@@ 1 Introduction
Motivated by this question, [GKVZ22] develop a theoretical framework to understand the power and limitations of such undetectable backdoors. [GKVZ22] prove that under standard cryptographic assumptions it is impossible to detect the existence of backdoors when we only have black-box access to the ML model. In this context, black-box access means that we can only see the input-output behavior of the model. We provide a more detailed comparison with [GKVZ22] in Section 1.2.
Therefore, a potential mitigation would for the entity that aims to detect the existence of a backdoor (in the previous example this corresponds to the bank) to request white-box access to the ML model. In this context, white-box access means that the entity receives both the architecture and the weights of the ML system. [GKVZ22] show that in some restricted cases, i.e., for random Fourier features [RR07], planting undetectable backdoors is possible even when the entity that tries to detect the backdoors has white-box access. Nevertheless, [GKVZ22] leave open the question of whether undetectability is possible for general models under white-box access.
Data Privacy & Obfuscation
A separate issue that arises with white-box access is that the details about the architecture and parameters of the ML models might reveal sensitive information, such as
  • Intellectual Property (IP): With white-box access to the system someone can reverse-engineer and understand the underlying algorithms and logic used to train which compromises the intellectual property of the entity that produces the ML models.
  • Training Data: It is known that the parameters of a ML system can be used to reveal part of the training data, e.g., [SRS17]. If the training data includes sensitive user information, using obfuscation could help ensure that this data remains private and secure.
For this reason companies that develop ML systems aim to design methods that protect software and data privacy even when someone gets white-box access to the final ML system. Towards this goal, obfuscation is a very powerful tool that is applied for similar security reasons in a diverse set of computer science applications [SKK+16]. Roughly speaking, obfuscation is a procedure that gets a program as input and outputs another program, the obfuscated program, that should satisfy three desiderata [Bar02]: (i) it must have the same functionality (i.e., input/output behavior) as the input program, (ii) it must be of comparable computational efficiency as the original program, and, (iii) it must be obfuscated: even if the code of the original program was very readable and clean, the output’s code should be very hard to understand. We refer to [BGI+01, Bar02] and Section 5 for further discussion on why obfuscation is an important security tool against IP and data privacy attacks.
Motivated by this, we operate under the assumption that the training of the ML models follow the “honest obfuscated pipeline”. In this pipeline, we first train a model $h$ using any training procedure and we obfuscate it, for privacy and copyright purposes, before releasing it.
Honest Obfuscated Pipeline
training data $\to$ Train $\rightarrow$ ML model $h$rightarrow Obfuscation $\to$ obfuscated ML model $\widetilde{h}$
@@
In this work we develop a framework to understand the power and limitations of backdoor attacks with white-box access when the ML models are produced via the honest obfuscated pipeline. We operate under the assumption that the obfuscation step is implemented based on the celebrated cryptographic technique called indistinguishability obfuscation (iO) [BGI+01, JLS21]. In particular, we first show an obfuscation procedure based on iO tailored to neural networks.
Our main result is a general provably efficient construction of a backdoor for deep neural networks (DNNs) that is undetectable even when we have white-box access to the model, assuming that the obfuscation is implemented based on iO. Based on this general construction we also develop a technique for introducing backdoors even to language models (LMs).
Together with the results of [GKVZ22], our constructions show the importance of cryptographic techniques to better understand some fundamental risks of modern Machine Learning systems.
@@ 1.1 Our Results
In this section we give a high-level description of our main results. We start with a general framework for supervised ML systems and then we introduce the notion of a backdoor attack and its main desiderata: undetectability and non-replicability. Finally, we provide an informal statement of our results.
@@
In this work, we focus on obfuscated models. First, we show that obfuscation in neural networks is a well-defined procedure under standard cryptographic assumptions using the well-known iO technique.
Theorem 1.1 (Obfuscation for Neural Networks).
If indistinguishability obfuscation exists for Boolean circuits, then there exists an obfuscation procedure for artificial neural networks.
This result is based on the existence of a transformation from Boolean circuits to ANNs and vice versa, formally introduced in Section 3.2. The procedure of Theorem 1.1 and, hence its proof, is explicitly presented in Section 4.1 and Remark 4.1.
Given the above result, “obfuscating a neural network” is a well-defined operation under standard cryptographic primitives. Hence, we can now provide our working assumption.
Assumption 1.2 (Honest Obfuscated Pipeline).
The training pipeline is defined as follows:
  1. 1.
    We train a model using $\mathsf{Train}$ and obtain a neural network classifier $h=\textnormal{sgn}(f)$ [^1] For simplicity, we assume that the neural network $f$ is a mapping from $[0,1]^{n}\to[0,1]$. Hence, we define $\textnormal{sgn}(x)\triangleq\mathbb{1}\{2x-1>0\}$ for $x\in[0,1]$..
  2. 2.
    Then, we obfuscate the neural network $f$ using the procedure of Theorem 1.1 to get $\widetilde{f}$.
  3. 3.
    Finally, we output the obfuscated neural network classifier $\widetilde{h}=\textnormal{sgn}(\widetilde{f}).$
@@
This leads us to the following definitions of undetectability and non-replicability (both introduced by [GKVZ22]) that a strong backdoor attack should satisfy. For short, we will write $\widetilde{h}\sim\mathsf{Backdoor}$ to denote a backdoored model
Definition 1.3 (Undetectability [GKVZ22]; Informal, see Definition 2.3).
We will say that a backdoor $(\mathsf{Backdoor},\mathsf{Activate})$ is undetectable with respect to the training procedure $\mathsf{Train}$ if for any data distribution $\mathcal{D}$, it is impossible to efficiently distinguish between $h$ and $\widetilde{h}$, where $h\sim\mathsf{Train}$ and $\widetilde{h}\sim\mathsf{Backdoor}$.
@@
    The backdoor is called white-box undetectable if it is impossible to efficiently distinguish between $h$ and $\widetilde{h}$ even with white-box access to $h$ and $\widetilde{h}$ (we receive a complete explicit description of the trained models, e.g., model’s architecture and weights).
@@
    The backdoor is called black-box undetectable if it is impossible to efficiently distinguish between $h$ and $\widetilde{h}$ when we only receive black-box query access to the trained models.
Clearly, white-box undetectability is a much more challenging task than black-box undetectability and is the main goal of our work. Black-box undetectability is by now very well understood based on the results of [GKVZ22], see also Table 1.
@@
If we assume that one-way functions and indistinguishability obfuscation exist, then for every honest obfuscated pipeline (satisfying 1.2) there exists a backdoor attack $(\mathsf{Backdoor},\mathsf{Activate})$ for ANNs that is both white-box undetectable and non-replicable.
As observed in Table 1, we know that black-box undetectable and non-replicable backdoors can be injected to arbitrary training procedures [GKVZ22]. However, this is unlikely for white-box undetectable ones. Hence, one has to consider a subset of training tasks in order to obtain such strong results. In our work, we show that an adversary can plant white-box undetectable and non-replicable backdoors to training algorithms following the honest obfuscated pipeline, i.e., an arbitrary training method followed by an obfuscation step. Prior to our result, only well-structured training processes, namely the RFF method, was known to admit a white-box undetectable backdoor [GKVZ22].
@@
Training Process Undetectability Non-Replicability
[GKVZ22] Arbitrary Black-Box Yes
[GKVZ22] RFF White-Box No
Our Work Obfuscated Pipeline White-Box Yes
Table 1: Comparison with Prior Work.
@@
In order to obtain the backdoor attack of Theorem 1.5 we develop a set of tools appearing in Section 3. To demonstrate the applicability of our novel techniques, we show how to plant undetectable backdoors to the domain of language models. This problem has been raised in various surveys such as [HCSS21, ASR+24] and has been experimentally investigated in a sequence of works e.g., in [KJTC23, XJX+24, WCP+23, ZWT+23, ZJT+24, RT23, RCM+24, HDM+24, ZZJW21]. As a first step, we introduce the notion of backdoor attacks in language models (see Definition 4.7). Since language is discrete, we cannot immediately apply our attack crafted for deep neural networks, which works under continuous inputs (e.g., by modifying the least significant input bits). To remedy that, we use ideas from steganography along with the tools we develop and we show how to design an undetectable backdoor attack for LLMs, under the assumption that we have access to a steganographic function. We refer to Section 4.2 for details.
@@
Conclusion & Open Questions
Given the plethora of applications of Machine Learning in general, and neural networks in particular, questions regarding the trustworthiness of publicly released models naturally arise. In particular, before deploying a neural network we need to guarantee that no backdoors have been injected allowing bad actors to arbitrarily control the model behavior. In this paper, we investigate the existence of backdoor attacks to obfuscated neural networks which are undetectable even when given white-box access. The notion of obfuscation that we consider is the well-studied and mathematically founded indistinguishability obfuscation (iO). We also show how our techniques can inspire backdoor schemes in large language models when combined with ideas from steganography.
While our constructions are purely theoretical, we leave as an interesting direction how to use heuristic obfuscation methods to show practical instantiations of our constructions. Another interesting open question is whether cryptographic schemes weaker than iO suffice to show backdoor undetectability in the white-box model.
@@ 1.2 Related Work
The work of [GKVZ22] is the closest to our work. At a high level, they provide two sets of results. Their first result is a black-box undetectable backdoor. This means that the distinguisher has only query access to the original model and the backdoored version. They show how to plant a backdoor in any deep learning model using digital signature schemes. Their construction guarantees that, given only query access, it is computationally infeasible, under standard cryptographic assumptions, to find even a single input where the original model and the backdoored one differ. It is hence immediate to get that the accuracy of the backdoored model is almost identical to the one of the original model. Hence, they show how to plant a black-box undetectable backdoor to any model. Their backdoor is also non-replicable. This result appears in the first row of Table 1. Our result applies to the more general scenario of white-box undetectability and hence is not comparable.
@@
The second set of results in [GKVZ22] is about planting white-box undetectable backdoors for specific algorithms (hence, they do not apply to all deep learning models, but very specific ones). The main model that their white-box attacks apply to is the RFF model of [RR07]. See also the second row of Table 1. Let us examine how [GKVZ22] add backdoors that are white-box undetectable. They first commit to a parameterized model (in particular, the Random Fourier Features (RFF) model of [RR07] or a random 1-layer ReLU NN), and then the honest algorithm commits to a random initialization procedure (e.g., every weight is sampled from $\mathcal{N}(0,I)$). After that, the backdoor algorithm samples the initialization of the model from an “adversarial” distribution that is industinguishable from the committed honest distribution and then uses the committed train procedure (e.g., executes the RFF algorithm faithfully on the given training data). Their main result is that, essentially, they can plant a backdoor in RFF that is white-box undetectable under the hardness of the Continuous Learning with Errors (CLWE) problem of [BRST21]. Our result aims to achieve further generality: we show that any training procedure followed by an obfuscation step can be backdoored in a white-box and non-replicable manner (see the third row of Table 1).
@@
The work of [MMS21] is similar to our work in terms of techniques but their goal is different: they show how to produce a model that (i) perfectly fits the training data, (ii) misclassifies everything else, and, (iii) is indistinguishable from one that generalizes well. At a technical level, [MMS21] also use indistinguishability obfuscation and signature schemes. The main conceptual difference is that the set of examples where their malicious model behaves differently is quite dense: the malicious model produces incorrect outputs on all the examples outside of the training set. In our setting and that of [GKVZ22], the changes in the model’s behavior are essentially measure zero on the population level and a backdoored model generalizes exactly as the original model.
[HCK22] study what they call “handcrafted” backdoors, to distinguish from prior works that focus exclusively on data poisoning. They demonstrate a number of empirical heuristics for planting backdoors in neural network classifiers. [GJMM20] show that there are learning tasks and associated classifiers, which are robust to adversarial examples, but only to a computationally-bounded adversaries. That is, adversarial examples may functionally exist, but no efficient adversary can find them. Their construction is similar to the black-box planting of [GKVZ22]. A different notion of backdoors has been extensively studied in the data poisoning literature [MB21, KLM+23, JHO24, HKSO21, TLM18, CLL+17, GLDGG19]. In this case, one wants to modify some part of the training data (and their labels) to plant a backdoor in the final classifier, without tampering with any other part of the training process. See also [SJI+22] for some connections between backdoors attacks and transfer learning. On the other side, there are various works studying backdoor detection [ASSK23].
The line of work on adversarial examples [IST+19, AEIK18, SZS+13] is also relevant to backdoors. Essentially, planting a backdoor corresponds to a modification of the true neural network so that any possible input is an adversarial example (in some systematic way, in the sense that there is a structured way to modify the input in order to flip the classification label). Various applied and theoretical works study the notion of adversarial robustness, which is also relevant to our work [RSL18, WK18, SNG+19, BLPR19]. Finally, backdoors have been extensively studied in cryptography. [YY97] formalized cryptographic backdoors and discussed ways that cryptographic techniques can themselves be used to insert backdoors in cryptographic systems. This approach is very similar to both [GKVZ22] and our work on how to use cryptographic tool to inject backdoors in deep learning models.
@@ 2.2 Planting Backdoors
Formally we give the following definition of a backdoor attack that consists of two algorithms $\mathsf{Backdoor}$ and $\mathsf{Activate}$.
@@
  • $\mathsf{Backdoor}$ is a probabilistic polynomial-time training algorithm that has oracle access to a distribution $\mathcal{D}$ over $\mathcal{X}\times\mathcal{Y}$ and outputs an ANN $\widetilde{f}:\mathcal{X}\to\mathbb{R}$ and a backdoor key $\mathsf{bk}$. The corresponding classifier is defined by $\widetilde{h}(x)=\textnormal{sgn}(\widetilde{f}(x))$.
@@
In general, we will either write $(\widetilde{h},\mathsf{bk})\sim\mathsf{Backdoor}$ or simply $\widetilde{h}\sim\mathsf{Backdoor}$ if the backdoor key $\mathsf{bk}$ is not crucial for the discussion.
@@ 2.3 Backdoor Detection
Definition 2.3 (Backdoor Detection [GKVZ22]).
We say that a model backdoor $(\mathsf{Backdoor},\mathsf{Activate})$, as in Definition 2.2, is white-box (resp. black-box) undetectable with respect to the training procedure $\mathsf{Train}$ if it satisfies that $h\sim\mathsf{Train}$ and $\widetilde{h}\sim\mathsf{Backdoor}$ are white-box (resp. black-box) undetectable in the sense that the two induced distributions are computationally-indistinguishable by probabilistic polynomial-time algorithms (see Definition 2.1) even if one has access to the complete description (architecture and weights) of the output neural networks (resp. if one has black-box query access to the output neural networks) .
In summary, for white-box undetectability, we ask whether there exists an efficient procedure that can be used to “hide” planted backdoors in neural networks in a very strong sense: even if one observes the output neural network’s architecture and weights, they cannot efficiently detect whether a backdoor was injected or not.